# Software Framework for Optically Accessible Quantum Memory Using Group-IV Color Centers in Diamond

### v2

@@ V Discussion and Outlook
Our executable quantum memory component leverages tin-vacancy (G4V) centers to store photonic time-bin qubits and demonstrates both robustness and versatility. We optimize $\pi/2$ pulse sequences using either microwave or optical control and tune cavity parameters to simulate the spin–photon entanglement that underlies the read-in and read-out operations. From these simulations we derive the corresponding Kraus-operators, thereby achieving the component’s central objective. The full module is available in [46]. Users may select optical or microwave control for both read-in and read-out, and in all cases they can configure photon-generation fidelity, bandwidth and the nanophotonic system’s temperature. If microwave control is chosen, strain as well as DC and AC magnetic field strengths and orientations are also adjustable; under optical control these parameters remain fixed.
The module can be further extended to enhance its versatility. For example, lifting the current constraints on optical control to permit arbitrary magnetic field strengths would broaden the accessible parameter space. Other group-IV vacancy centers, such as the silicon, germanium, and lead vacancies [62], can also be included, serving an even broader research community.
@@ Acknowledgements
Author Contributions
Y.S. and G.P. conceptualized the research, with Y.S. conducting the simulations and data analysis. M.B. performed the simulations specific to microwave spin control. G.P. and T.S. developed the core idea and provided overall project supervision. All authors contributed to writing and refining the manuscript.
@@
Group-IV color centers (G4V) in diamond have four states $\{|k\rangle\}$ $k\in G=\{1,2,3,4\}$ in the ground state manifold and four states $\{|l\rangle\}$ $l\in E=\{A,B,C,D\}$ in the excited state manifold. Therefore, the Hamiltonian of the G4V coupled to a single cavity mode is given by
$H=\sum_{m=2}^{8}\epsilon_{m}|m\rangle\langle m|+\omega_{c}a_{c}^{\dagger}a_{c}+\left(\sum_{k\in G,l\in E}g_{kl}\sigma_{kl}+g_{kl}^{*}\sigma_{lk}\right)\left(a+a^{\dagger}\right)$ (31)
@@
where $\alpha=1,2$ and $\beta=A,B$, are the G4V natural decay rates based on Fermi’s golden rule [47]. Here, $\alpha=1/137$ is the fine structure constant, $e$ is the elementary charge. The G4V natural decay rates relate to the decay time $T_{1}$ as $\gamma=1/T_{1}$. For the SiV (SnV), $T_{1}=1.7(4.5)$ ns [65, 47], with the Debye-Waller factor $0.8(0.6)$ [1, 47]. Additionally, we define the theoretical cooperativities [66] as a function of the coupling strength $g_{kl}$, full width at half maximum of the spectral curve $\gamma_{kl}$, and full width at half maximum of the spectral curve $2\kappa$ as
@@
$\rho_{+}=\int{}_{\omega}\langle+|\psi\rangle\langle\psi|+\rangle_{\omega}\,{\rm d}\omega,$ (70)
$\rho_{-}=\int{}_{\omega}\langle-|\psi\rangle\langle\psi|-\rangle_{\omega}\,{\rm d}\omega.$ (71)
The respective entries are
@@
as the target state. We define $\eta_{\rm sp}={\rm tr}(\tilde{\rho}_{\rm sp})$ as the success probability of spin-photon entanglement and $\rho_{\rm sp}=\tilde{\rho}_{\rm sp}/\eta_{\rm sp}$ as the state. We optimize for cavity parameters $(\kappa,\omega_{c})$ and emission central frequency $\omega_{0}$ such that the fidelity
$F_{\rm sp}=\langle\rm Bell|\rho_{\rm sp}|\rm Bell\rangle$ (82)
with $|\rm Bell\rangle=\frac{1}{\sqrt{2}}(|1\rangle+|2\rangle)$ is maximized for $\alpha=\beta=1/\sqrt{2}$.
@@
For the read-in process we model a photon source emitting broadband photons of bandwidth $\gamma$ and fidelity $F_{\rm ph}$ and assume an imperfect $\pi/2$ spin rotation produced by either microwave or optical spin control. Starting point is the photonic qubit
$|\psi_{0}\rangle=\frac{1}{\sqrt{2}}(|E\rangle+|L\rangle).$ (83)
To model imperfect photon pair generation we apply a depolarizing channel [68], i.e.
$\rho=(1-\epsilon)\rho_{0}+\epsilon I/2$ (84)
with $\epsilon=2(1-F_{\rm ph})$, the fidelity $F_{\rm ph}$ of the photon pair generated by the quantum dot and the pure state $\rho_{0}=|\psi_{0}\rangle\langle\psi_{0}|$. For simplicity we employ the notation
$\rho=\sum\rho_{IJKM}|IJ\rangle\langle KM|$ (85)
@@
$\sum\rho_{IK}|I\rangle\langle K|1\rangle\langle 1|=\sum\rho_{IK}a^{I}\bar{a}^{K}|1\rangle\langle 1|$ (87)
$\xrightarrow[]{\text{early reflection}}$ $\sum\rho_{IK}(\delta_{IE}\mathcal{D}_{1}(a^{E})+\delta_{IL}a^{L})$ (88)
$(\delta_{KE}\bar{\mathcal{D}}_{1}(a^{E})+\delta_{KL}\bar{a}^{L})|1\rangle\langle 1|$ (89)
$\xrightarrow[]{\pi/2\,\,\text{rotation}}$ $\sum\rho_{IK}\sum\Lambda^{(mk)}(\delta_{IE}\mathcal{D}_{1}(a^{E})+\delta_{IL}a^{L})$ (90)
$(\delta_{KE}\bar{\mathcal{D}}_{1}(a^{E})+\delta_{KL}\bar{a}^{L})|m\rangle\langle k|$ (91)
$\xrightarrow[]{\text{late reflection}}$ $\sum\rho_{IK}\sum\Lambda^{(mk)}(\delta_{IE}\mathcal{D}_{1}(a^{E})+\delta_{IL}\mathcal{D}_{m}(a^{L}))$ (92)
$(\delta_{KE}\bar{\mathcal{D}}_{1}(a^{E})+\delta_{KL}\bar{\mathcal{D}}_{k}(\bar{a}^{L}))|m\rangle\langle k|$ (93)
where $\mathcal{D}_{k}(a)$ denotes the outcoming mode $a_{\rm out}$ from the time evolution according to the Heisenberg-Langevin equations shown in Eqs. (55)-(64) initializing $\sigma_{kk}(0)=1$ for an input mode $a_{\rm in}$. After the $X$-measurement the state is
$\rho_{\pm}=\frac{1}{2}\sum_{IK}\sum_{m,k}\Lambda^{(mk)}\left(\delta_{IE}\delta_{KE}I_{1}\pm\delta_{IE}\delta_{KL}I_{k}\pm\delta_{IL}\delta_{KE}I_{2m-1}+\delta_{IL}\delta_{KL}I_{2(m-1)+k}\right)$ (94)
@@
with $\mathcal{N}=\gamma/e_{0}^{2}$ and $T$ sufficiently large.
@@
For the read-out process we assume a perfect CPHASE gate between the spin and photon and no depolarization during photon generation for simplicity. The process requires an extra single photon source emitting single photons of the form $|\psi_{\rm ph}\rangle=1/\sqrt{2}(|e\rangle+|l\rangle)$ with the corresponding density matrix $\rho_{\rm ph}=|\psi_{\rm ph}\rangle\langle\psi_{\rm ph}|$. For the computation of Kraus-operators for the read-out process we apply the reflection scheme on the full basis of $2\times 2$ matrices for the spin. The initial state is $\rho^{ij}=\rho_{\rm ph}\otimes|i\rangle\langle j|$. The controlled phase gates applied here are $U_{e}={\rm diag}(-1,1,1,1)$ and $U_{l}={\rm diag}(1,1,-1,1)$. We assume an imperfect $\pi/2$ rotation for the spin qubit driven by either optical or microwave control modeled by the Kraus-operators $\{K_{m}\}$. After the late reflection the state reads
$\rho_{\rm lr}^{ij}=U_{l}\left(\sum_{m=1}^{4}K_{m}\left(U_{e}\rho^{ij}U_{e}^{\dagger}\right)K_{m}^{\dagger}\right)U_{l}^{\dagger},\quad i,j=1,2$ (98)
The measurement outcomes in the $Z$-basis are $\rho_{11}^{ij}=\langle 1|\rho^{ij}|1\rangle$ and $\rho_{22}^{ij}=\langle 2|\rho^{ij}|2\rangle$.
Evaluating the photonic qubit for $i,j=1,2$ results in the Kraus-operators for the read-out process.
APPENDIX E KRAUS-OPERATORS
E.1 Optical Control
$K_{1}^{\rm in}$ $=\begin{pmatrix}0.527&0.527\\ 0.385+0.015\,{\rm i}&-0.379-0.013\,{\rm i}\end{pmatrix}$ (99)
$K_{2}^{\rm in}$ $=\begin{pmatrix}0.152-0.005\,{\rm i}&0.152-0.005\,{\rm i}\\ -0.204-0.001\,{\rm i}&0.215\end{pmatrix}$ (100)
$K_{3}^{\rm in}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (101)
$K_{4}^{\rm in}$ $=\begin{pmatrix}-0.001&-0.001\\ 0.08&0.078\end{pmatrix}$ (102)
$K_{1}^{\rm out}$ $=\begin{pmatrix}0.51&0.417-0.022\,{\rm i}\\ 0.51&-0.417+0.022\,{\rm i}\end{pmatrix}$ (103)
$K_{2}^{\rm out}$ $=\begin{pmatrix}-0.205-0.011\,{\rm i}&0.251\\ -0.205-0.011\,{\rm i}&-0.251\end{pmatrix}$ (104)
$K_{3}^{\rm out}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (105)
$K_{4}^{\rm out}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (106)
E.2 Microwave Control
$K_{1}^{\rm in}$ $=\begin{pmatrix}0.446&0.446\\ 0.442+0.024\,{\rm i}&-0.301+0.244\,{\rm i}\end{pmatrix}$ (107)
$K_{2}^{\rm in}$ $=\begin{pmatrix}0.045+0.037\,{\rm i}&0.045+0.037\,{\rm i}\\ 0.050+0.044\,{\rm i}&0.210\end{pmatrix}$ (108)
$K_{3}^{\rm in}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (109)
$K_{4}^{\rm in}$ $=\begin{pmatrix}-0.027+0.001\,{\rm i}&-0.027+0.001\,{\rm i}\\ 0.053+0\,{\rm i}&-0.002+0.001\,{\rm i}\end{pmatrix}$ (110)
$K_{1}^{\rm out}$ $=\begin{pmatrix}-0.499-0.024\,{\rm i}&-0.500\\ -0.499-0.024\,{\rm i}&0.500\end{pmatrix}$ (111)
$K_{2}^{\rm out}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (112)
$K_{3}^{\rm out}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (113)
$K_{4}^{\rm out}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (114)

### v3

@@ V Discussion and Outlook
Our executable quantum memory component leverages negatively charged silicon- and tin-vacancy (G4V) centers to store photonic time-bin qubits and demonstrates both robustness and versatility. The module computes the Kraus operators describing the read-in and read-out processes by internally optimizing all relevant control pulses and cavity parameters, including the locally optimal optical controls from [6]. The full implementation is available in [49].
Users may select optical control for the SnV- or microwave control for the SiV- and SnV- for both read-in and read-out, and in all cases they can configure photon-generation fidelity of the auxiliary photon source, bandwidth of the incoming photon for read-in and read-out and the nanophotonic system’s temperature. The user can also choose between the optimized phase gate and manually chosen parameters to produce a phase gate as well as the optimized spin gate and manually set propagated spin states. If microwave control is chosen, strain as well as DC and AC magnetic field strengths and orientations are also adjustable; under optical control these parameters remain fixed.
The module can be further extended to enhance its versatility. For example, lifting the current constraints on optical control to permit arbitrary magnetic field strengths would broaden the accessible parameter space. Other group-IV vacancy centers, such as the germanium, and lead vacancies [73], can also be included, serving an even broader research community. In future iterations of the software module we plan to include the carbon nuclear spin as the quantum memory to increase the coherence time.
@@ Acknowledgements
Group-IV color centers (G4V) in diamond have four states $\{|k\rangle\}$ with $k\in G=\{1,2,3,4\}$ in the ground state manifold and four states $\{|l\rangle\}$ with $l\in E=\{A,B,C,D\}$ in the excited state manifold. Therefore, the Hamiltonian of the G4V coupled to a single cavity mode is given by
$H=\sum_{m\in G\cup E}\epsilon_{m}|m\rangle\langle m|+\omega_{c}a_{c}^{\dagger}a_{c}+\left(\sum_{k\in G,l\in E}g_{kl}\sigma_{kl}+g_{kl}^{*}\sigma_{lk}\right)\left(a+a^{\dagger}\right)$ (36)
@@
where $\alpha=1,2$ and $\beta=A,B$, are the G4V natural decay rates based on Fermi’s golden rule [50]. Here, $\alpha=1/137$ is the fine structure constant, $e$ is the elementary charge. The G4V natural decay rates relate to the decay time $T_{1}$ as $\gamma=1/T_{1}$. For the SiV-(SnV), $T_{1}=1.7(4.5)$ ns [76, 50], with the Debye-Waller factor $0.8(0.6)$ [1, 50]. Additionally, we define the theoretical cooperativities [77] as a function of the coupling strength $g_{kl}$, full width at half maximum of the spectral curve $\gamma_{kl}$, and full width at half maximum of the spectral curve $2\kappa$ as
@@
$\rho_{+}=\int_{\mathbb{R}}{}_{\omega}\langle+|\psi\rangle\langle\psi|+\rangle_{\omega}\,{\rm d}\omega,$ (75)
$\rho_{-}=\int_{\mathbb{R}}{}_{\omega}\langle-|\psi\rangle\langle\psi|-\rangle_{\omega}\,{\rm d}\omega$ (76)
with $|+\rangle_{\omega}=\frac{|\omega\rangle_{E}+|\omega\rangle_{L}}{\sqrt{2}}$. The respective entries are
@@
as the target state. We define $\eta={\rm tr}(\rho)$ as the success probability of spin-photon entanglement. We optimize for cavity parameters $(\kappa,\omega_{c})$ and emission central frequency $\omega_{0}$ such that the fidelity
$F=\frac{1}{\eta}\langle\phi|\rho|\phi\rangle$ (87)
with $|\phi\rangle=\frac{1}{\sqrt{2}}(|1\rangle+|2\rangle)$ is maximized for $\alpha=\beta=1/\sqrt{2}$.
@@
For the read-in process we model a photon source emitting broadband photons of bandwidth $\gamma$ and assume an imperfect $\pi/2$ spin rotation produced by either microwave or optical spin control. In the following we use superscripts as matrix elements an use capital letters for indexing the photonic qubit $(I,K)$ and a lowercase letter $(m,k)$ for the spin qubit index. To model the read-in process with Kraus-operators the reflection scheme is applied to each basis state, i.e.
$\rho_{\rm ph}^{(IK)}=|I\rangle\langle K|$ (88)
@@
$\rho_{0}^{(IK)}:=|I\rangle\langle K|1\rangle\langle 1|=a^{I}\bar{a}^{K}|1\rangle\langle 1|$ (90)
$\xrightarrow[]{\text{early reflection}}$ $(\delta_{IE}\mathcal{D}_{1}(a^{E})+\delta_{IL}a^{L})(\delta_{KE}\bar{\mathcal{D}}_{1}(a^{E})+\delta_{KL}\bar{a}^{L})|1\rangle\langle 1|$ (91)
$\xrightarrow[]{\pi/2\,\,\text{rotation}}$ $\sum_{m,k}\Lambda^{(mk)}(\delta_{IE}\mathcal{D}_{1}(a^{E})+\delta_{IL}a^{L})(\delta_{KE}\bar{\mathcal{D}}_{1}(a^{E})+\delta_{KL}\bar{a}^{L})|m\rangle\langle k|$ (92)
$\xrightarrow[]{\text{late reflection}}$ $\sum_{m,k}\Lambda^{(mk)}(\delta_{IE}\mathcal{D}_{1}(a^{E})+\delta_{IL}\mathcal{D}_{m}(a^{L}))(\delta_{KE}\bar{\mathcal{D}}_{1}(a^{E})+\delta_{KL}\bar{\mathcal{D}}_{k}(a^{L}))|m\rangle\langle k|=:\rho_{\rm sp-ph}^{(IK)}$ (93)
where $I,K=E,L$ and $\mathcal{D}_{k}(a)$ denotes the output mode $a_{\rm out}$ from the time evolution according to the Heisenberg-Langevin equations shown in Eqs. (60)-(69) initializing $\sigma_{kk}(0)=1$ for an input mode $a_{\rm in}$. We measure in the $X$-basis which has the basis states $|\pm\rangle=\frac{|E\rangle\pm|L\rangle}{\sqrt{2}}$. After the $X$-measurement the state is
$\rho_{\pm}^{(IK)}=\langle\pm|\rho_{\rm sp-ph}^{(IK)}|\pm\rangle=\frac{1}{2}\sum_{m,k}\Lambda^{(mk)}\left(\delta_{IE}\delta_{KE}I_{1}\pm\delta_{IE}\delta_{KL}I_{k}\pm\delta_{IL}\delta_{KE}I_{2m-1}+\delta_{IL}\delta_{KL}I_{2(m-1)+k}\right)|m\rangle\langle k|$ (94)
@@
where $\mathcal{N}=\gamma/e_{0}^{2}$ and $T$ sufficiently large. We compute Kraus-operators for the measurement in $|+\rangle$ and $|$-$\rangle$ separately using the Choi-Jamiolkowski-Isomorphism [69, 70, 51].
@@
To model the read-out process we calculate Kraus-operators which map the spin state to the photonic state up to a rotation. To achieve that aim we assume that the auxiliary photon source emits a photon with the fidelity $F_{\rm aux}$. The ideal state is $|\psi_{\rm ph}\rangle=1/\sqrt{2}(|E\rangle+|L\rangle)$. The depolarized photonic qubit is [79]
$\rho_{\rm ph}=(1-\epsilon)|\psi_{\rm ph}\rangle\langle\psi_{\rm ph}|+\epsilon I/2,\quad\epsilon=2(1-F_{\rm aux}).$ (98)
We perform the reflection scheme on the spin-photon state $\rho_{0}=\rho_{\rm ph}\otimes|m\rangle\langle n|$ for $m,n=1,2$ and perform a $Z$-measurement on the spin qubit. We assume that the $\pi/2$ rotated spin state of $|m\rangle\langle n|$ is given by $\mathcal{D}_{\pi/2}(|m\rangle\langle n|)=\sum_{p,q=1}^{2}\Lambda_{mn}^{(pq)}|p\rangle\langle q|$. The reflection scheme reads
$\rho_{0}=\sum_{I,K}\rho_{\rm ph}^{(IK)}a^{I}\bar{a}^{K}|m\rangle\langle n|$ (99)
$\xrightarrow[]{\text{early reflection}}\sum_{I,K}\rho_{\rm ph}^{(IK)}(\delta_{IE}\mathcal{D}_{m}(a^{E})+\delta_{IL}a^{L})(\delta_{KE}\bar{D}_{n}(a^{K})+\delta_{KL}\bar{a}^{L})|m\rangle\langle n|$ (100)
$\xrightarrow[]{\pi/2\,\,\text{rotation}}\sum_{I,K}\sum_{m,n}\rho_{\rm ph}^{(IK)}(\delta_{IE}\mathcal{D}_{m}(a^{E})+\delta_{IL}a^{L})(\delta_{KE}\bar{D}_{n}(a^{K})+\delta_{KL}\bar{a}^{L})\Lambda_{mn}^{pq}|p\rangle\langle q|$ (101)
$\xrightarrow[]{\text{late reflection}}\sum_{I,K}\sum_{m,n}\rho_{\rm ph}^{(IK)}(\delta_{IE}\mathcal{D}_{m}(a^{E})+\delta_{IL}\mathcal{D}_{p}(a^{L}))(\delta_{KE}\bar{D}_{n}(a^{K})+\delta_{KL}\bar{\mathcal{D}_{q}}(a^{L})\Lambda_{mn}^{pq}|p\rangle\langle q|=:\rho_{\rm sp-ph}$ (102)
We perform a $Z$-measurement of the spin qubit. It is
$\rho_{1}=\langle 1|\rho_{\rm sp-ph}|1\rangle=\sum_{I,K}\rho_{\rm ph}^{(IK)}(\delta_{IE}\mathcal{D}_{m}(a^{E})+\delta_{IL}\mathcal{D}_{1}(a^{L}))(\delta_{KE}\bar{D}_{n}(a^{K})+\delta_{KL}\bar{\mathcal{D}_{1}}(a^{L})\Lambda_{mn}^{(11)},$ (103)
$\rho_{2}=\langle 2|\rho_{\rm sp-ph}|2\rangle=\sum_{I,K}\rho_{\rm ph}^{(IK)}(\delta_{IE}\mathcal{D}_{m}(a^{E})+\delta_{IL}\mathcal{D}_{2}(a^{L}))(\delta_{KE}\bar{D}_{n}(a^{K})+\delta_{KL}\bar{\mathcal{D}_{2}}(a^{L})\Lambda_{mn}^{(22)}.$ (104)
To evaluate the components of the photonic qubits $\rho_{1}$ and $\rho_{2}$ in the time-bin basis $\{|E\rangle,|L\rangle\}$ we evaluate the state components, i.e.
$\langle E|\rho_{i}|E\rangle=\mathcal{N}\int_{0}^{T}\Lambda_{mn}^{(ii)}\rho_{\rm ph}^{(EE)}\mathcal{D}_{m}(a_{\rm in})\bar{\mathcal{D}}_{n}(a_{\rm in})\,{\rm d}t=\mathcal{N}\Lambda_{mn}^{(ii)}\rho_{\rm ph}^{(EE)}I_{2(m-1)+n},$ (105)
$\langle E|\rho_{i}|L\rangle=\mathcal{N}\int_{0}^{T}\Lambda_{mn}^{(ii)}\rho_{\rm ph}^{(EL)}\mathcal{D}_{m}(a_{\rm in})\bar{\mathcal{D}}_{i}(a_{\rm in})\,{\rm d}t=\mathcal{N}\Lambda_{mn}^{(ii)}\rho_{\rm ph}^{(EL)}I_{2(m-1)+i},$ (106)
$\langle L|\rho_{i}|E\rangle=\mathcal{N}\int_{0}^{T}\Lambda_{mn}^{(ii)}\rho_{\rm ph}^{(LE)}\mathcal{D}_{i}(a_{\rm in})\bar{\mathcal{D}}_{n}(a_{\rm in})\,{\rm d}t=\mathcal{N}\Lambda_{mn}^{(ii)}\rho_{\rm ph}^{(LE)}I_{2(i-1)+n},$ (107)
$\langle L|\rho_{i}|L\rangle=\mathcal{N}\int_{0}^{T}\Lambda_{mn}^{(ii)}\rho_{\rm ph}^{(LL)}|\mathcal{D}_{i}(a_{\rm in})|^{2}\,{\rm d}t=\mathcal{N}\Lambda_{mn}^{(ii)}\rho_{\rm ph}^{LL}I_{3(i-1)+1}$ (108)
for $i=1,2$ where I is shown in the Appendix C. We compute Kraus-operators for the measurement in $|1\rangle$ and $|2\rangle$ separately using the Choi-Jamiolkowski-Isomorphism [69, 70, 51].
APPENDIX E QUANTUM MEMORY PERFORMANCE
To estimate the quantum memory performance we evaluate the average trace and infidelity for the read-in and read-out process of a photonic qubit. We evaluate the following steps:
  1. 1.
    Compute Kraus-operators $\{K_{i}^{+},K_{i}^{-}\}$ for the read-in process of any photonic qubit
  2. 2.
    Compute Kraus-operators $\{K_{i}^{1},K_{i}^{2}\}$ for the read-out process of the spin qubit
  3. 3.
    Consider the photonic qubit
    $|\psi_{\rm ph}(\theta,\phi)\rangle=\cos(\theta/2)|E\rangle+e^{{\rm i}\phi}\sin(\theta/2)|L\rangle,\quad\theta\in[0,\pi],\phi\in[0,2\pi)$ (109)
    with the respective density matrix $\rho_{\rm ph}(\theta,\phi)=|\psi_{\rm ph}(\theta,\phi)\rangle\langle\psi_{\rm ph}(\theta,\phi)|$. We evaluate the spin states after the read-in process, i.e.
    $\rho_{+}(\theta,\phi)=\sum_{i=1}^{4}K_{i}^{+}\rho_{\rm ph}(\theta,\phi)(K_{i}^{+})^{\dagger},$ (110)
    $\rho_{-}(\theta,\phi)=\sum_{i=1}^{4}K_{i}^{-}\rho_{\rm ph}(\theta,\phi)(K_{i}^{-})^{\dagger}.$ (111)
    Using the auxiliary photon source we use the reflection scheme to entangle the spin states $\rho_{+}$ and $\rho_{-}$ with the auxiliary photon. After the read-out process the photonic states are
    $\rho_{+,1}(\theta,\phi)=\sum_{i=1}^{4}K_{i}^{1}\rho_{\rm+}(\theta,\phi)(K_{i}^{1})^{\dagger},$ (112)
    $\rho_{+,2}(\theta,\phi)=\sum_{i=1}^{4}K_{i}^{2}\rho_{\rm+}(\theta,\phi)(K_{i}^{2})^{\dagger},$ (113)
    $\rho_{-,1}(\theta,\phi)=\sum_{i=1}^{4}K_{i}^{1}\rho_{\rm-}(\theta,\phi)(K_{i}^{1})^{\dagger},$ (114)
    $\rho_{-,2}(\theta,\phi)=\sum_{i=1}^{4}K_{i}^{2}\rho_{\rm-}(\theta,\phi)(K_{i}^{2})^{\dagger}.$ (115)
    The total photonic states are
    $\rho_{\rm ph,+}(\theta,\phi)=\rho_{+,1}(\theta,\phi)+\sigma_{z}\sigma_{x}\rho_{+,2}(\theta,\phi)\sigma_{x}\sigma_{z},$ (116)
    $\rho_{\rm ph,-}(\theta,\phi)=\rho_{-,1}(\theta,\phi)+\sigma_{z}\sigma_{x}\rho_{-,2}(\theta,\phi)\sigma_{x}\sigma_{z}$ (117)
    with $\sigma_{x}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$. From these states the total photonic state after read-out is given by
    $\tilde{\rho}_{\rm ph}(\theta,\phi)=\rho_{\rm ph,+}(\theta,\phi)+\sigma_{z}\rho_{\rm ph,-}(\theta,\phi)\sigma_{z}$ (118)
    The trace of the photon after the joint read-in and read-out process is subsequently given by ${\rm tr}(\tilde{\rho}_{\rm ph}(\theta,\phi))$ and the fidelity is $F(\theta,\phi)=\frac{1}{\eta(\theta,\phi)}\langle\psi_{\rm ph}(\theta,\phi)|\tilde{\rho}_{\rm ph}(\theta,\phi)|\psi_{\rm ph}(\theta,\phi)\rangle$.
  4. 4.
    Compute the Bloch average for the fidelity and trace
    $\langle F\rangle:=\frac{1}{4\pi}\int_{0}^{\pi}\int_{0}^{2\pi}F(\theta,\phi)\sin(\theta)\,{\rm d}\theta\,{\rm d}\phi,\quad\langle{\rm tr}(\rho_{\rm ph})\rangle:=\frac{1}{4\pi}\int_{0}^{\pi}\int_{0}^{2\pi}{\rm tr}(\tilde{\rho}_{\rm ph}(\theta,\phi))\sin(\theta)\,{\rm d}\theta\,{\rm d}\phi.$ (119)
APPENDIX F Supporting Data
Table 1: Microwave powers and total processing times for qubit storage and retrieval $(P^{\rm MW}/\mu\text{W},T_{1}/{\rm ns})$ assuming a fiber length $L=100$ m, a detector’s dead time $T_{m}=100$ ps, the refractive index of diamond $n=2.417$, the bandwidth of the photon for read-in and read-out $\gamma_{\rm in,out}=0.1$ GHz, for $\theta_{\rm dc}=\pi/2$, $\theta_{\rm ac}=0$ with the field strength $B_{\rm ac}=1.0$ $(3.7)$ mT for the SiV-(SnV-)as quantum memory
Center $B_{\rm dc\perp}$ (T) $P^{\rm MW}$ $(\mu\text{W})$ $T_{1}$ (ns)
SiV- $0.3$ $359$ $509$
SiV- $0.4$ $202$ $509$
SnV- $0.3$ $876$ $531$
SnV- $0.4$ $493$ $531$
Table 2: Microwave powers and total processing times for qubit storage and retrieval $(P^{\rm MW}/\mu\text{W},T_{1}/{\rm ns})$ assuming a fiber length $L=100$ m, a detector’s dead time $T_{m}=100$ ps, the refractive index of diamond $n=2.417$, the bandwidth of the photon for read-in and read-out $\gamma_{\rm in,out}=0.1$ GHz, for $\theta_{\rm dc}=0$, $\theta_{\rm ac}=\pi/2$ with the field strength $B_{\rm dc}=0.3$ T and $B_{\rm ac}=1.0$ $(3.7)$ mT for the SiV- (SnV-) as quantum memory
$P^{\rm MW}$ $(\mu\text{W})$ $T_{1}$ $({\rm ns})$
SiV- $274$ $539$
SnV- $16$ $2802$
Table 3: Laser power saturation ratio and total processing times for qubit storage and retrieval $(s_{1},s_{2},T_{1}/{\rm ns})$ assuming a fiber length $L=100$ m, a detector’s dead time $T_{m}=100$ ps, the refractive index of diamond $n=2.417$, the bandwidth of the photon for read-in and read-out $\gamma_{\rm in,out}=0.1$ GHz.
$B_{\rm dc}$ (T) $0.3$ $1.0$ $3.0$
$s_{1}$ $0.430$ $0.744$ $0.093$
$s_{2}$ $0.042$ $0.066$ $0.004$
$T_{1}$ (ns) $501.29$ $500.55$ $503.00$